Nate Johnston

Potomac Falls, VA, USA · iam@natejohnston.info · (703) 473-4079

Work Experience

Red Hat Remote Engineering Manager 7/2020 - present

• Managed two teams: first a team dedicated to enhancing advanced networking services within OpenStack like Load Balancer as a Service and DNS as a Service, then a team tasked with optimizing architecture and tooling for highly performant NFV use cases of OpenStack.

• Managed a team of 14 associates across 3 continents to implement scrum for the first time, significantly increasing the predictability of team delivery. Led the team in multiple technological innovations, ensuring those innovations built support within the company and became part of the product offering.

• Represented Red Hat Engineering on a number of high profile customer support escalations, providing calm communication of progress and managed expectations with diplomacy.

• Led a cross-functional Process Improvement team, personally designed the Jira workflow for the OpenStack product. Achieved consensus around my proposals through advocacy and outreach, with significant efficiencies across the product teams as a result.

• visibility check

Red Hat Remote Principal Product Developer 6/2018 - 6/2020

• Contributed code and participated as a core maintainer of the OpenStack Neutron project, implementing features and remediating defects for layer 2 and layer 3 networking within OpenStack. My biggest achievement was implementing efficiencies around bulk creation of resources to provide significant speed and performance improvements.

• Elected to the OpenStack Technical Committee, the governing board for the OpenStack project. Formulated a proposal, that was accepted, to provide projects within the OpenStack umbrella a dispersed project maintainer role as opposed to the traditional approach of having a single primary developer.

Comcast Reston, VA Principal Product Developer 1/2010 - 10/2011; 10/2013 - 5/2015; 10/2016 - 6/2018

• Team lead and software developer programming primarily in Go and Ruby, working to engineer a holistic datacenter automation solution for Comcast.

• Responsible for creating technical designs, code reviews, and ensuring code and operational quality in an Agile/Scrum devops environment.

• Developed processes to extract and analyze utilization and billing data from AWS, Azure, and OpenStack.

• Helped build a Spinnaker-as-a-Service offering to offer hosted Spinnaker at enterprise scale.

• Designed, deployed, and provided full lifecycle support for containerized applications running in both Kubernetes and Mesos.

• Replaced the HP Server Automation service with multiple off the shelf open source systems.

• Designed, deployed, and supported a large multi-tenant Puppet Master infrastructure.

Comcast Reston, VA Principal OpenStack Developer 6/2015 - 10/2016

• Core Contributor for both OpenStack neutron and neutron-fwaas.

• In the Mitaka cycle, I was part of a team of developers across multiple companies that collaborated to implement DSCP management within the Quality of Service (QoS) framework in neutron. I also presented about DSCP at the Austin 2016 OpenStack summit.

• In the Newton cycle, I led the neutron-fwaas team to implement FWaaS v2, a significant new version of the project. I coordinated heavily with the neutron team, and contributed heavily to both neutron and neutron-fwaas.

Comcast Reston, VA Principal Systems Architect 11/2011 - 9/2013

• Developed and deployed an innovative server quality assurance and auditing system to ensure consistent delivery of quality server resources.

• Established a more rigorous data model for the operational CMDB that tracked server inventory and assignments, to ensure data quality and clarity of relationships between data sources.

AOL Reston, VA/Dulles, VA Principal System Administrator 6/2001 - 12/2009

• SRE before it had a name. System administration, in-depth application administration, automation development, creating new monitoring, metrics, and CMDB infrastructure. Leader in cross-enterprise task forces.

Skills

Programming Languages:  Go, Python, Ruby, Shell
Container Technologies:  Docker, Kubernetes, Mesos, Marathon, Consul, Vault
Clouds:  OpenStack, AWS, Azure
Configuration Management: Puppet, Ansible, Terraform

Education

University of Virginia Charlottesville, VA BA History 1993 - 1997